OpenStudy (anonymous):

Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius. I'm not sure how to find the arc degree...

OpenStudy (campbell_st):

 the triangle formed by the chord and radii is equilateral (all sides 8 inches). So the angle subtended at the centre is 60 degrees the problems is now Area of the sector - area of the triangle. find the perpendicular height of the triangle  by pythagoras the height is 8^2 - 4^2 h\[h = \sqrt{48} = 4\sqrt{3}\] area of the triangle is \[A = \frac{1}{2} \times 8 \times 4\sqrt{3} = 16\sqrt{3}\] next the sector the angle at the centre is 60 so the fraction of the circle is 60/360 or 1/6 Area of the sector is \[A = \frac{1}{6} \times \pi \times 8^2\] then the area of the segment is \[A = \frac{1}{6} \pi 8^2 - 16\sqrt{3}\] I'll leave you to evaluate the answer